HETAV PATEL

Grade 11,

in cross product why i*i=j*j=k*k=o. in dot product why i*i=j*j=k*k=1

in cross product why i*i=j*j=k*k=o.
in dot product why i*i=j*j=k*k=1

Grade:11

2 Answers

PASUPULETI GURU MAHESH
8 years ago
IN  CROSS PRODUCT A X B=|A|.|B| SIN ∂ (∂ INDICATE THETA)
IN DOT PRODUCT A.B=|A||B| COS∂
THEN ANGLE BETWEEN  I×I;J×J;K×K=0  DEGREES.
IN DOT PRODUCT COS(0)=1   THEN  I×I;J×J;K×K=1
IN  CROSS PRODUCT SIN(0)=0   THEN  I×I;J×J;K×K=0
 
SREEKANTH
7 years ago
IN CROSS PRODUCT I*I=I *I*SINA  WHERE A IS AQNGLE BETWEEN THE TWO UNIQUE VECTORS I I
                                      =I*I SIN(0) THE ANGLE BETWEEN TWO VECTORS IS ZERO
                                      =0
LIKE J*J=K*K=0
IN DOT PRODUCT I*I=I*I COS A  WHERE A IS AN ANGLE BETWEEN TWO VECTORS I I
                                 =I*I COS(0)    THE ANGLE BETWEEN TWO PARALLEL VECTORS IS ZERO
                                 =I^2    THE MAGNITUDE OF UNIT VECTOR IS 1
                                 =1
LIKE J*J=K*K=1